OpenStudy (anonymous):

What is the second term in the expansion of (2x + 7)^6 ?

OpenStudy (anonymous):

A. 1344x^4 B. 1344x^5 C. 23,520x^4 D. 23,520x^5

OpenStudy (anonymous):

lol i think i must be doing the same course as u cause this is everything im going lol

geerky42 (geerky42):

\[\Large(a+b)^n = \sum_{k=0}^n\left(\begin{matrix}n \\ k\end{matrix}\right)x^{n-k}y^k\]Second term is when k = 1 so \(\left(\begin{matrix}6 \\ 1\end{matrix}\right)(2x)^{6-1}7^1 = 6\cdot2^5\cdot 7x^5 = \boxed{1344x^5}\)

OpenStudy (anonymous):

Thanks!!

geerky42 (geerky42):

You're welcome.